If 10mL of 0.3 M KOH are required to neutralize 30mL of gastric juice (HCl), what is the molarity of the gastric juice
1
Expert's answer
2020-05-08T14:12:00-0400

C1V1=C2V2;

C2=C1V1/V2=0.3*10/30=0.1 M

Answer: 0.1 M